OpenStudy (anonymous):

Water is leaking out of an inverted conical tank at a rate of 9800 cubic centimeters per min at the same time that water is being pumped into the tank at a constant rate. The tank has height 8 meters and the diameter at the top is 3 meters. If the water level is rising at a rate of 16 centimeters per minute when the height of the water is 5 meters, find the rate at which water is being pumped into the tank in cubic centimeters per minute

OpenStudy (anonymous):

V' in = V' ( currently left ) + V' leak V = ( π /3) r² h = ( π /3) ( 1.5/ 8)² h³ -> V' = ( 1.5/ 8)² π h² h' plug h = 5, h' = .16 m/min : ... @katleynn782 Now the rest is just calculating. Don't forget convert into cm³ !